# How do Massive Primordial Black Holes Impact the Formation of the First Stars and Galaxies?

### v1

@@ 1 Introduction
The emerging JWST first-galaxy phenomenology is providing a powerful stress test for standard $\Lambda$CDM cosmology (e.g., Inayoshi et al., 2022; Boylan-Kolchin, 2023). To alleviate the tension and account for this rapid galaxy assembly and the highly efficient star formation within a short cosmic time ($\sim 300~{}\rm Myr$), multiple solutions have been proposed, including starburst regimes that are relatively free from supernova and other stellar feedback (see e.g., Dekel et al., 2023; Shen et al., 2023; Boylan-Kolchin, 2024) and boosted light-to-mass ratios from a top-heavy initial mass function or peculiar stellar evolution (e.g., Trinca et al., 2024; Wang et al., 2024; Liu et al., 2024b). Additionally, modifications to the underlying cosmology have also been suggested, such as accelerated structure formation by invoking a blue tilt in the primordial power spectrum (e.g., Padmanabhan & Loeb, 2023; Hirano & Yoshida, 2024) or early dark energy (e.g., Shen et al., 2024). In extreme cases, such scenarios could trigger first star formation as early as $z\gtrsim 100$ (Hirano et al., 2015; Ito & Omukai, 2024).
Furthermore, JWST has revealed an abundant population of SMBHs at $z\gtrsim 8$ (e.g., Goulding et al., 2023; Larson et al., 2023; Bogdán et al., 2024; Greene et al., 2024; Kovács et al., 2024; Maiolino et al., 2024a; Natarajan et al., 2024). Among them is the extremely distant SMBH discovered in the well-constrained GN-z11 system (Tacchella et al., 2023; Maiolino et al., 2024b). At slightly lower redshifts ($z=4-9$) lie a previously unknown class of objects, ultra-compact and dust-obscured sources often referred to as “little red dots” (LRDs) that may represent an early evolutionary phase of black hole growth. LRDs therefore may provide unique insights into the formation and evolution of early cosmic structures (e.g., Labbé et al., 2023; Labbe et al., 2023; Kokorev et al., 2024; Kocevski et al., 2024; Leung et al., 2024; Taylor et al., 2024). Intriguingly, most of the newly-discovered high-$z$ SMBHs exhibit ratios of their inferred mass to the host stellar mass exceeding the local relation (e.g., Pacucci et al., 2023; Inayoshi & Ichikawa, 2024). This suggests that the conditions under which these early structures formed were significantly different from those of galaxies in the present-day Universe.
@@
Another promising scenario invokes massive primordial black holes (PBHs) as seeds for early SMBHs (for analytical work see, e.g., Düchting, 2004; Bernal et al., 2018; De Luca et al., 2023; Dayal, 2024). PBHs as a well-motivated dark matter candidate (reviewed in Carr & Kühnel, 2020), theorized to form from the collapse of overdensities shortly after the Big Bang (Zel’dovich & Novikov, 1967; Hawking, 1971; Carr, 1975; Belotsky et al., 2019; Escrivà, 2022), present a promising mechanism for addressing some of the challenges posed by recent JWST observations (for a review of various PBH formation scenarios, see Escrivà et al., 2024). Unlike stellar remnants or DCBHs, PBHs do not require a pre-existing star-forming environment, making them an attractive candidate for explaining the presence of SMBHs at very high redshifts (Yuan et al., 2024; Huang et al., 2024; Ziparo et al., 2024). Even if PBHs were to constitute only a small fraction of the dark matter, they could significantly influence cosmic history by generating a rich phenomenology (for a summary of the constraints, see Carr et al., 2021b). Moreover, there is now an increasing set of observations which might be allocated to PBHs, such as various microlensing events (e.g., Niikura et al., 2019; Mediavilla et al., 2017; Wyrzykowski & Mandel, 2020; Hawkins, 2022, 2024), correlations in the cosmic x-ray and infrared background fluctuations (Kashlinsky et al., 2005), absence of certain ultrafaint dwarf galaxies (e.g., Clesse & García-Bellido, 2018), several LIGO/Virgo/KAGRA subsolar black hole merger candidates (Phukon et al., 2021; LVK Collaboration, 2023; Morrás et al., 2023), or certain of supernovæ (Calcium-rich gap-transients) which do not trace the stellar population but could well be explained by PBH-triggered white dwarf explosion (Smirnov et al., 2024). These and further hints for PBHs are discussed in Carr et al. (2024).
@@
However, much of the previous PBH research relied heavily on analytical models or dark matter-only simulations, which do not capture the complex interactions between PBHs, surrounding gas, and the larger-scale environment. Black hole feedback—whereby radiation and winds from accreting black holes influence their surroundings—can alter the thermodynamic state of the interstellar medium (e.g., Lu et al., 2021; Takhistov et al., 2022; Casanueva-Villarreal et al., 2024). This feedback, if sufficiently strong, may even suppress the collapse of gas into stars [^1] One caveat is that, since PBHs are also able to accelerate structure formation, more minihalos would form earlier. When considering both effects together, there may be no significant change to the overall Pop III star formation rate (Liu et al., 2022)., thereby delaying star formation, as previously explored in simulations with stellar mass ($10-100\,\mathrm{M}_{\odot}$) PBHs (Liu et al., 2022). Moreover, analytical models are inherently limited in their ability to simulate the three-dimensional complexity of accretion disks around black holes and the subsequent feedback processes. Such limitations restrict our understanding of the non-linear, multi-scale interactions that are crucial for accurately modeling the evolution of early black hole seeds and the first galaxies, thus emphasizing the need for hydrodynamical simulations.
@@ 2 Methodology
Table 1: Summary of key simulation parameters and results. $L$ is the length of the simulation box in comoving kpc. $z_{\rm ini}$ is the initial redshift where the simulation starts, and $z_{*}$ the redshift when stars begin to form around the central PBH. $N_{\rm eff}$ is the total number of particles within the simulation box. $\epsilon_{r}$ is the thermal feedback coupling efficiency. $\Delta m_{\rm BH}$ is the total mass accreted by the central black hole by the final redshift, and $m_{\star}$ denotes the total mass of the stars that formed in the PBH-hosting halo by the end of the simulation ($z=9$). STR is a flag indicating whether the relative streaming of PDM and gas particles is included (✓) or not (✗), and the value represents the amplitude with respect to the root-mean-square streaming velocity $\sigma_{\rm b\chi}$. BH_LW is another flag to control whether we include (✓) the local LW feedback from BH accretion or not (✗).
@@ 2.1 Pathfinder Simulation and Initial Conditions
We start with a PDM-only pathfinder simulation, denoted as PDMonly, using the initial conditions generated by the MUSIC code (Hahn & Abel, 2011) at $z_{\rm eq}=3400$, which marks the beginning of the matter-dominated epoch and of the growth of structure, placing a single PBH at the center of the simulation box. We use $m_{\rm BH}$ to represent the PBH mass, initially set to $10^{6}\,\mathrm{M}_{\odot}$ throughout all simulation runs that contain the PBH. We assign a box side length of $L\sim 250\ \rm ckpc$ throughout, with $N_{\rm PDM}=256^{3}$ setting the resolution of PDM particles. To add perturbations and calculate the velocities of PDM particles surrounding the PBH, we adopt the numerical recipe from previous work (Ali-Haïmoud & Kamionkowski, 2017; Inman & Ali-Haïmoud, 2019b; Liu & Bromm, 2023; Zhang et al., 2024a), employing the Zel’dovich approximation (Zel’dovich, 1970).
@@ 3.1 PBH Accretion and Feedback
This phase of efficient accretion is quickly terminated by shock waves generated from the accretion feedback, which raise the gas temperature near the PBH and inhibit further gas collapse. Notably, when the feedback efficiency is higher, the accretion burst is weaker because stronger shocks are launched more rapidly. Across all simulation runs, this large variations of the accretion rate are observed, driven by the competition between cold gas inflow from the intergalactic medium and the thermal feedback from BH accretion.
@@ 4 Limitations and Caveats
Due to limitations in our computational resources and the scale of the simulation, our models stop at $z=9$, which restricts our ability to study the long-term evolution of PBH-stellar systems and their host halos at lower redshifts. In addition, at $z\lesssim 10$, thermal feedback from PBH accretion begins to heat up the entire simulation box, rendering the evolution of star clusters around PBH-seeded halos less physical. This excessive heating slows down gas cooling, reduces the star formation rate, and leads to a more extended stellar density distribution than would be expected in a fully resolved system. Future simulations with improved resolution and larger volumes will be necessary to better capture the long-term effects of PBH accretion feedback on star formation and galaxy evolution, particularly at $z\lesssim 10$. Furthermore, the limited box size prevents us from accounting for large-scale structures that may form around a PBH-seeded halo. If a PBH is embedded in an overdensity peak whose scale far exceeds that of the PBH-seeded halo, we would expect a merger with a nearby galaxy, which could boost the stellar mass, disrupt accretion, and alter the $m_{\mathrm{BH}}/m_{\star}$ and $m_{\mathrm{BH}}/m_{h}$ ratios. Therefore, while the current work is valid for a galaxy seeded by an isolated PBH, additional simulations will be needed to explain the early formation of galaxies with $m_{\star}\gg 10^{6}\,\mathrm{M}_{\odot}$ observed at $z\gtrsim 10$ (e.g., Finkelstein et al., 2022; Castellano et al., 2022; Naidu et al., 2022; Castellano et al., 2023; Labbé et al., 2023; Adams et al., 2024; Donnan et al., 2024).
@@ 5 Summary and Conclusions
  • PBHs exhibit a dual influence on the timing and location of star formation. On the one hand, PBHs accelerate structure formation by seeding dark matter halos and gravitationally attracting gas to form dense regions. On the other hand, radiative and thermal feedback from accretion can delay gas cooling, suppressing star formation in cases of strong feedback. In cases with weaker feedback, gas cooling proceeds efficiently, leading to higher densities and enhanced star formation. This dual behavior highlights the complexity of PBH effects, which can both suppress and promote star formation depending on the feedback strength.
@@
  • Under strong Lyman–Werner (LW) feedback, the star formation process is quenched as the dense gas in the vicinity of the PBH cannot efficiently cool, resulting in near-isothermal evolution dominated by atomic hydrogen cooling.
@@
To distinguish PBHs from other SMBH formation pathways, further observations of early AGN with frontier facilities, such as the JWST or the Roman Space Telescope (RST), will be crucial to achieve statistically significant samples of the overall SMBH demographics at high redshifts. Furthermore, the dense star clusters in the vicinity of a PBH, as can be inferred from the stellar density profile in Fig. 6, could imply a high frequency of tidal disruption events, which may be observable with JWST or in future surveys with RST (Inayoshi et al., 2024).
The additional detection of high-$z$ AGN in radio-wavebands would allow a more complete physical characterization, when cross-correlating with the near-infrared (JWST) results (e.g., Ibar et al., 2008). Future missions, such as the Square Kilometer Array (SKA) (for constraints on massive PBHs, see Bernal et al., 2018), the Hydrogen Epoch of Reionization Array (HERA) (Koopmans et al., 2021), and lunar-based 21-cm cosmology observatories (Burns et al., 2021), offer unique opportunities to probe PBH-driven processes and provide deeper physical insight. For instance, these facilities could detect thermal or radiative signatures originating from PBHs during the cosmic dark ages ($z\sim 30-200$), providing indirect evidence of their influence on the early IGM and cosmic radiation background (see also, e.g., Kashlinsky, 2016; Hasinger, 2020). Additionally, such missions could explore the extremely early formation of stars, which might arise due to the blue tilt of the power spectrum introduced by PBHs (see also Hirano et al., 2015).
@@
We would like to acknowledge fruitful discussions with Steve Finkelstein and Seiji Fujimoto, our colleagues in UT’s Cosmic Frontier Center, as well as with Hui Li, Kohei Inoyashi, Haibo Yu, and James Gurian. The authors acknowledge the Texas Advanced Computing Center (TACC) for providing HPC resources under allocation AST23026. BL gratefully acknowledges the funding of the Royal Society University Research Fellowship and the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC 2181/1 - 390900948 (the Heidelberg STRUCTURES Excellence Cluster). MBK acknowledges support from NNSF grants AST-1910346, AST-2108962, and AST-2408247; NASA grant 80NSSC22K0827; HST-GO-16686, HST-AR-17028, HST-AR-17043, JWST-GO-03788, and JWST-AR-06278 from the Space Telescope Science Institute, which is operated by AURA, Inc., under NASA contract NAS5-26555; and from the Samuel T. and Fern Yanagisawa Regents Professorship in Astronomy at UT Austin. Here, we first demonstrate the effect of relative DM-baryon streaming on the onset of first star formation (Fig. 10), where an ultra-early collapse occurs in the vicinity of the PBH. On the other hand, Fig. 11 illustrates the formation of star clusters around the PBH under conditions of reduced feedback strength. In this case, the PBH-seeded galaxy develops a disk-like configuration in both its gaseous and stellar components.

### v2

@@ 1 Introduction
The emerging JWST first-galaxy phenomenology is providing a powerful stress test for standard $\Lambda$CDM cosmology (e.g., Inayoshi et al., 2022; Boylan-Kolchin, 2023). To alleviate the tension and account for this rapid galaxy assembly and the highly efficient star formation within a short cosmic time ($\sim 300~{}\rm Myr$), multiple solutions have been proposed, including starburst regimes that are relatively free from supernova and other stellar feedback (see e.g., Dekel et al., 2023; Shen et al., 2023; Boylan-Kolchin, 2025) and boosted light-to-mass ratios from a top-heavy initial mass function or peculiar stellar evolution (e.g., Trinca et al., 2024; Wang et al., 2024; Liu et al., 2024b). Additionally, modifications to the underlying cosmology have also been suggested, such as accelerated structure formation by invoking a blue tilt in the primordial power spectrum (e.g., Padmanabhan & Loeb, 2023; Hirano & Yoshida, 2024) or early dark energy (e.g., Shen et al., 2024). In extreme cases, such scenarios could trigger first star formation as early as $z\gtrsim 100$ (Hirano et al., 2015; Ito & Omukai, 2024).
Furthermore, JWST has revealed an abundant population of SMBHs at $z\gtrsim 8$ (e.g., Goulding et al., 2023; Larson et al., 2023; Bogdán et al., 2024; Greene et al., 2024; Kovács et al., 2024; Maiolino et al., 2024a; Natarajan et al., 2024; Napolitano et al., 2024). Among them is the extremely distant SMBH discovered in the well-constrained GN-z11 system (Tacchella et al., 2023; Maiolino et al., 2024b). At slightly lower redshifts ($z=4-9$) lie a previously unknown class of objects, ultra-compact and dust-obscured sources often referred to as “little red dots” (LRDs) that may represent an early evolutionary phase of black hole growth. LRDs therefore may provide unique insight into the formation and evolution of early cosmic structures (e.g., Labbé et al., 2023; Kokorev et al., 2024; Kocevski et al., 2024; Leung et al., 2024; Taylor et al., 2024; Labbe et al., 2025). Intriguingly, most of the newly-discovered high-$z$ SMBHs exhibit ratios of their inferred mass to the host stellar mass exceeding the local relation (e.g., Pacucci et al., 2023; Inayoshi & Ichikawa, 2024). This suggests that the conditions under which these early structures formed were significantly different from those of galaxies in the present-day Universe.
@@
Another promising scenario invokes massive primordial black holes (PBHs) as seeds for early SMBHs (for analytical work see, e.g., Düchting, 2004; Bernal et al., 2018; De Luca et al., 2023; Dayal, 2024). PBHs as a well-motivated dark matter candidate (reviewed in Carr & Kühnel, 2020), theorized to form from the collapse of overdensities shortly after the Big Bang (Zel’dovich & Novikov, 1967; Hawking, 1971; Carr, 1975; Belotsky et al., 2019; Escrivà, 2022), present a plausible mechanism for addressing some of the challenges posed by recent JWST observations (for a review of various PBH formation scenarios, see Escrivà et al., 2024). Unlike stellar remnants or DCBHs, PBHs do not require a pre-existing star-forming environment, making them an attractive candidate for explaining the presence of SMBHs at very high redshifts (Yuan et al., 2024; Huang et al., 2024; Ziparo et al., 2025).
Even if PBHs were to constitute only a small fraction of the dark matter, they could significantly influence cosmic history by generating a rich phenomenology (for a summary of the constraints, see Carr et al., 2021b). Moreover, there is now an increasing set of observations which might be allocated to PBHs, such as various microlensing events (e.g., Niikura et al., 2019; Mediavilla et al., 2017; Wyrzykowski & Mandel, 2020; Hawkins, 2022, 2024), correlations in the cosmic x-ray and infrared background fluctuations (Kashlinsky et al., 2005), absence of certain ultra-faint dwarf galaxies (e.g., Clesse & García-Bellido, 2018), several LIGO/Virgo/KAGRA subsolar black hole merger candidates (Phukon et al., 2021; LVK Collaboration, 2023; Morrás et al., 2023), or select supernovæ (Calcium-rich gap-transients) which do not trace the stellar population, but could well be explained by PBH-triggered white dwarf explosions (Smirnov et al., 2024). These and further hints for PBHs are discussed in Carr et al. (2024).
@@
However, much of the previous PBH research relied heavily on analytical models or dark-matter-only simulations, which do not capture the complex interactions between PBHs, surrounding gas, and the larger-scale environment. Black hole feedback —- where radiation and winds from accreting black holes influence their surroundings —- can alter the thermodynamic state of the interstellar medium (e.g., Lu et al., 2021; Takhistov et al., 2022; Casanueva-Villarreal et al., 2024). This feedback, if sufficiently strong, may even suppress the collapse of gas into stars [^1] One caveat is that, since PBHs are also able to accelerate structure formation, more minihalos would form earlier. When considering both effects together, there may be no significant change to the overall Pop III star formation rate (Liu et al., 2022)., thereby delaying star formation, as previously explored in simulations with stellar mass ($10-100\,\mathrm{M}_{\odot}$) PBHs (Liu et al., 2022). Moreover, analytical models are inherently limited in their ability to simulate the three-dimensional complexity of accretion disks around black holes and the subsequent feedback processes. Such limitations restrict our understanding of the non-linear, multi-scale interactions that are crucial for accurately modeling the evolution of early black hole seeds and the first galaxies, thus emphasizing the need for hydrodynamical simulations.
@@ 2 Methodology
Table 1: Summary of key simulation parameters and results. $L$ is the length of the simulation box in comoving kpc. $z_{\rm ini}$ is the initial redshift where the simulation starts, and $z_{*}$ the redshift when stars begin to form around the central PBH. $N_{\rm eff}$ is the total number of particles within the simulation box. $\epsilon_{r}$ is the thermal feedback coupling efficiency. $\Delta m_{\rm BH}$ is the total mass accreted by the central black hole by the end of the simulation (note that some runs terminate at $z>9$), and $m_{\star}$ is the total stellar mass formed within the PBH-hosting halo, or within the PDM halo in the case of the CDM run, by $z=9$. STR is a flag indicating whether the relative streaming of PDM and gas particles is included (✓) or not (✗), and the value represents the amplitude with respect to the root-mean-square streaming velocity $\sigma_{\rm b\chi}$. BH_LW is another flag to control whether we include (✓) the local LW feedback from BH accretion or not (✗).
@@ 2.1 Pathfinder Simulation and Initial Conditions
We start with a PDM-only pathfinder simulation, denoted as PDMonly, using the initial conditions generated by the MUSIC code (Hahn & Abel, 2011) at $z_{\rm eq}=3400$, which marks the beginning of the matter-dominated epoch and the growth of structure, placing a single PBH at the center of the simulation box. We use $m_{\rm BH}$ to represent the PBH mass, initially set to $10^{6}\,\mathrm{M}_{\odot}$ throughout all simulation runs that contain the PBH [^4] In this work, we effectively assume that the mass fraction of PBHs in the PDM component is less than $6\times 10^{-4}$, representing an upper limit set by the ratio of the PBH mass to the total mass enclosed within the simulation volume (calculated for the average cosmological matter density). This value is consistent with existing constraints from large-scale structure formation (Carr & Silk, 2018) and dynamical friction effects (Carr & Sakellariadou, 1999). It also agrees with simulated results from previous work, exploring varying PBH density fractions (e.g., Zhang et al., 2024a; Casanueva-Villarreal et al., 2025).. We assign a box side length of $L\sim 250\ \rm ckpc$ throughout, with $N_{\rm PDM}=256^{3}$ setting the resolution of PDM particles. To add perturbations and calculate the velocities of PDM particles surrounding the PBH, we adopt the numerical recipe from previous work (Ali-Haïmoud & Kamionkowski, 2017; Inman & Ali-Haïmoud, 2019b; Liu & Bromm, 2023; Zhang et al., 2024a), employing the Zel’dovich approximation (Zel’dovich, 1970).
@@ 3.1 PBH Accretion and Feedback
This phase of efficient accretion is quickly terminated by shock waves generated from the accretion feedback, which raise the gas temperature near the PBH and inhibit further gas collapse. Notably, when the feedback efficiency is higher, the accretion burst is weaker because stronger shocks are launched more rapidly. Across all simulation runs, these large variations of the accretion rate are observed, driven by the competition between cold gas inflow from the intergalactic medium and the thermal feedback from BH accretion.
@@ 4 Limitations and Caveats
Due to limitations in our computational resources and the scale of the simulation, our models stop at $z=9$, restricting our ability to study the long-term evolution of PBH-stellar systems and their host halos at lower redshifts. In addition, at $z\lesssim 10$, thermal feedback from PBH accretion begins to heat up the entire simulation box, rendering the evolution of star clusters around PBH-seeded halos less physical. This excessive heating slows down gas cooling, reduces the star formation rate, and leads to a more extended stellar density distribution than would be expected in a fully resolved system. Future simulations with improved resolution and larger volumes will be necessary to better capture the long-term effects of PBH accretion feedback on star formation and galaxy evolution, particularly at $z\lesssim 10$. Furthermore, the limited box size prevents us from accounting for large-scale structures that may form around a PBH-seeded halo. If a PBH is embedded in an overdensity peak whose scale far exceeds that of the PBH-seeded halo, we would expect a merger with a nearby galaxy, which could boost the stellar mass, disrupt accretion, and alter the $m_{\mathrm{BH}}/m_{\star}$ and $m_{\mathrm{BH}}/m_{h}$ ratios. Therefore, while the current work is valid for a galaxy seeded by an isolated PBH, additional simulations will be needed to explain the early formation of galaxies with $m_{\star}\gg 10^{6}\,\mathrm{M}_{\odot}$ observed at $z\gtrsim 10$ (e.g., Finkelstein et al., 2022; Castellano et al., 2022; Naidu et al., 2022; Castellano et al., 2023; Labbé et al., 2023; Adams et al., 2024; Donnan et al., 2024).
@@ 5 Summary and Conclusions
  • PBHs exhibit a dual influence on the timing and location of star formation. On the one hand, PBHs accelerate structure formation by seeding dark matter halos and gravitationally attracting gas to form dense regions. On the other hand, radiative and thermal feedback from accretion can delay gas cooling, suppressing star formation in cases of strong feedback, and star formation also needs to compete with PBH accretion for the dense gas supply. In cases with weaker yet non-negligible feedback, gas cooling proceeds efficiently, leading to higher densities and enhanced star formation. However, under extremely weak feedback, the PBH devours all dense gas, thus preventing star formation. These diverse behaviors highlight the complexity of PBH effects, which can both suppress and promote star formation depending on the feedback strength.
@@
  • Under strong Lyman–Werner (LW) feedback, the star formation process is quenched (at least on the scales resolved by our simulation) as the dense gas in the vicinity of the PBH cannot efficiently cool, resulting in near-isothermal evolution dominated by atomic hydrogen cooling.
@@
To distinguish PBHs from other SMBH formation pathways, further observations of early AGN with frontier facilities, such as the JWST or the Roman Space Telescope (RST), will be crucial to achieve statistically significant samples of the overall SMBH demographics at high redshifts. Furthermore, the dense star clusters in the vicinity of a PBH, as can be inferred from the stellar density profile in Fig. 6, could imply a high frequency of tidal disruption events, which may be observable with JWST or in future surveys with RST (Inayoshi et al., 2024; Wang et al., 2025).
The additional detection of high-$z$ AGN in radio-wavebands (e.g., with the Atacama Large Millimeter/submillimeter Array, ALMA) would allow a more complete physical characterization, when cross-correlating with the near-infrared (JWST) results (e.g., Ibar et al., 2008; De Rossi & Bromm, 2023; Labbe et al., 2025). Future missions, such as the Square Kilometer Array (SKA) (for constraints on massive PBHs, see Bernal et al., 2018), the Hydrogen Epoch of Reionization Array (HERA) (Koopmans et al., 2021), and lunar-based 21-cm cosmology observatories (Burns et al., 2021), offer unique opportunities to probe PBH-driven processes and provide deeper physical insight. For instance, these facilities could detect thermal or radiative signatures originating from PBHs during the cosmic dark ages ($z\sim 30-200$), providing indirect evidence of their influence on the early IGM and cosmic radiation background (see also, e.g., Kashlinsky, 2016; Hasinger, 2020). Additionally, such missions could explore the extremely early formation of stars, which might arise due to the blue tilt of the power spectrum introduced by PBHs (see also Hirano et al., 2015; Hirano & Yoshida, 2024).
@@
We would like to acknowledge fruitful discussions with Steve Finkelstein and Seiji Fujimoto, our colleagues in UT’s Cosmic Frontier Center, as well as with Hui Li, Kohei Inoyashi, Haibo Yu, and James Gurian. The authors acknowledge the Texas Advanced Computing Center (TACC) for providing HPC resources under allocation AST23026. BL gratefully acknowledges the funding of the Royal Society University Research Fellowship and the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC 2181/1 - 390900948 (the Heidelberg STRUCTURES Excellence Cluster). MBK acknowledges support from NNSF grants AST-1910346, AST-2108962, and AST-2408247; NASA grant 80NSSC22K0827; HST-GO-16686, HST-AR-17028, HST-AR-17043, JWST-GO-03788, and JWST-AR-06278 from the Space Telescope Science Institute, which is operated by AURA, Inc., under NASA contract NAS5-26555; and from the Samuel T. and Fern Yanagisawa Regents Professorship in Astronomy at UT Austin. Here, we first demonstrate the effect of relative DM-baryon streaming on the onset of first star formation (Fig. 10), where an ultra-early collapse occurs in the vicinity of the PBH in the PBH_fd005_sstr simulation with $v_{\rm b\chi}=1.6~{}\sigma_{\rm b\chi}$. On the other hand, Fig. 11 illustrates the formation of star clusters around the PBH under conditions of reduced feedback strength in the PBH_fd0005 simulation with $\epsilon_{r}=0.05\%$. In this case, the PBH-seeded galaxy develops a disk-like configuration in both its gaseous and stellar components.